### v1

@@ 9 Ranking of the US states based on median incomes
Table 6: Average or expected absolute deviations for the Income Data
@@
We compare the frequentist and four Bayesian sets for $\theta$ by using their volume and length measures. We present these results in Table 5. Volumes of Bayesian sets are vastly smaller than the frequentist confidence sets. The HB elliptical set has the smallest volume. Average lengths of Bayesian sets are smaller than the average length of confidence set, again the HB elliptical set has the smallest average length. Here each of the volume numbers are scaled by a known constant $g$, and each of the length numbers are scaled by a known constant $h$.
In Table 6 we presented for each state average absolute deviations (cf. equation (14)) of the KWW ranks from the surrogate ranks based on 1989 income data from the census. We also presented the posterior expected absolute deviations, obtained from (14), for the credible solutions. The deviations for KWW method exceed the deviations for the HB method for all $51$ states, and those exceed $46$ deviations for the UB method excluding OK, GA, CO, NV, and DC. We find that the KWW method has the highest sum of average absolute deviations (799), about 344% larger than for the HB method, and about 116% larger than for the UB method. Clearly, the HB method is the best.
For the median income application we further explore comparison of credible distributions generated by various Bayesian models (UB and HB, with and without covariates) and methods (weighted and unweighted) among themselves and with the frequentist method of KWW in terms of average absolute deviations and posterior expected absolute deviations defined in equation (14). We plotted several of these posterior expected absolute deviations against the KWW average absolute deviations in Figure 5.
The left panel in the first row of Figure 5 is the plot using posterior expected absolute deviations from credible distributions from weighted method based on the HB model with covariates on the $y$-axis and the KWW deviations on the $x$-axis. On this plot we use the green circles for the elliptical method and the black stars for the rectangular method. This panel shows that all the points cluster below the $y=x$ line and closer to the $x$-axis, indicating that the credible distributions are vastly superior to the KWW method. It also shows that for the weighted method there are practically no differences between the deviations from credible distributions based on the rectangular or the elliptical credible sets for the mean vector $\theta$.
The right panel in the first row of that figure differs from the left panel only by use of no covariates in the HB model. We see that several points are above the $y=x$ line, indicating that the KWW solution for those states are better than the HB solution in the absence of covariates. Also, although many of these points that are below the $y=x$ line, a large number of points are above the line $y=5$ in contrast with the left panel where only about $15\%$ of the points are above the $y=5$ line, indicating that the credible distributions in the absence of good covariates are not as good as those created by using good covariates. These two panels convincingly demonstrate the importance of effective covariates for the ranking problem.
Figure 5: Various plots of expected or average absolute deviations
Refer to caption Refer to caption
With Covariates Covariates not used
Refer to caption Refer to caption
With Covariates No covariates, elliptical
Refer to caption Refer to caption
Covariates cannot be used Covariates cannot be used
The left panel on the second row mimics the left panel on the first row except that now the credible distributions are unweighted. The HB model employs covariates. As we have noticed in the first row, in this case also we see that the Bayesian solutions (both elliptical and rectangular) are vastly better than the KWW solution. However, while most of the points on the panel in the second row are above the line $y=5$, most are below that line in the panel on the first row. These two panels show the superiority of the weighted credible distributions. In the right panel on this row we plotted the deviations based on weighted, elliptical credible distributions for the UB model and the HB model without covariates against the KWW deviations. While overall the credible distributions show superiority over the KWW method, for about $10\%$ of the states the KWW method performed better. None of these methods utilized any covariates. This plot also shows that in the absence of a covariate the HB method loses its advantage over the UB method, the two credible distributions performed rather similarly.
In the last row of the figure, we compare various credible distributions (weighted vs. unweighted, elliptical vs. Cartesian) from the UB model against the confidence set provided by the KWW method in terms of posterior expected absolute deviations (resulting from the credible distributions) and average absolute deviations (resulting from the KWW method). While again the UB credible distributions show overall superiority over the KWW method, for about $10\%$ of the states the KWW method performed better. None of these methods utilized any covariates.
10 Summary and Conclusions
Inference for population means (or proportions) of subpopulations or subjects, is an important statistical problem. In survey sampling estimation of small area characteristics, in animal science and livestock research estimation of breeding values of various herds, in sports estimation of true performance of various players, and in healthcare estimation of treatment effectiveness for certain conditions treated by several health facilities, are some of the important problems in multi-population studies. The major focus in statistical inference for multiple populations has been estimation (point and interval) of means or proportions or similar measures from several subpopulations or individuals. In small area estimation where the main goal has been estimation of small area means, the companion problem of simultaneously estimating the ranks of the small areas is also of considerable importance but it has not gotten enough attention. Only Shen and Louis (1998) simultaneously considered estimation of small area means and ranks. Using the HB approach these authors considered point estimation of ranks but did not account for error in estimation of the rank vector for all small areas.
To account for estimation error for the rank vector of many subpopulations, in a pioneering article Klein et al. (2020) considered set estimation of overall ranking. Klein et al. (2020) created joint confidence set of the rank vector by employing a rectangular confidence set of the mean vector $\theta$. Their solution in equation (8) attributes $|\Lambda_{Li}|+1,\cdots,|\Lambda_{Li}|+|\Lambda_{Oi}|+1|$ as the possible values for the true rank ${\check{r}}_{i},i=1,\cdots,m.$ This solution, frequentist by construction, may be interpreted as a confidence distribution for ${\check{r}}_{i}$, uniformly distributed over the integers between the lower and the upper bounds presented above. As an alternative to this method, which cannot use auxiliary information or model the unknown true mean $\theta$, we introduced a Bayesian model for the mean vector. The model allows us to utilize useful auxiliary information as well as to develop various credible sets for $\theta$ including the HPD sets that are exact or nearly exact. We create a credible distribution for the true rank ${\check{r}}_{i}$, which is more informative than the confidence set by KWW. In our Bayesian proposal we can also allow more probable $\theta$’s in the credible set to have them bigger weights to the credible distribution.
Application of proposed credible distributions is computationally straightforward. We applied our solution to three applications, one to rank professional baseball players, and two to rank small areas based on unknown small area means. All three applications and a simulation study mimicking two of the applications demonstrated superior performance of the proposed method over the existing frequentist method. For the baseball example there are no reliable covariates but there are surrogate values of the true means from the players’ performance from the remainder of the season. Evaluation of frequentist confidence set and the credible distributions based on the average or expected absolute deviations of estimated ranks from the ranks of the surrogate values show better performance for the Bayesian solutions. In the absence of covariates, the UB credible distribution performs better than the HB credible distribution. For the commuting time example we neither have any covariates nor any representative values of the unknown true means. Even then in this example the credible distributions are “more conclusive”, “less dispersed” and “more illuminating” than the frequentist solution by Klein et al. (2020). For the median income estimation problem we have useful covariates as well as representative values (from the 1990 census) of the true means. Credible distributions from the HB model with the covariates performed the best when measured in terms average absolute deviations. Both credible distributions, from the UB model and the HB model, have better performance than the frequentist method. The HB credible distribution based on covariates is uniformly better than the other two solutions. However, the HB credible distribution without any covariates does not enjoy a uniform superiority over the other two solutions. These outcomes are confirmed by our simulation study for the setup of the baseball application. Availability of good covariates show the superiority of the HB credible distribution. Lack of covariates degrades the performance of the HB credible distribution, which is sometimes inferior to the UB credible distribution. In terms of size measures of the confidence or credible sets, the Bayesian solutions are vastly superior, and between the two Bayesian solutions, the HB-based credible sets are shorter or more “compact” than the UB-based credible sets.
Disclosure Statement
There are no competing interests to declare from the authors.
Funding
Appendix: size measures of elliptical sets
As an alternative to an orthotope (or hyperrectangular) credible set, where each component of $\theta$ varies freely, we have considered HPD sets which are approximately elliptical for our Bayesian models. Two size measures that appear to be appropriate for set estimators are their Lebesgue volumes and average lengths of the “sides” determining the set. These two measures are easily found for an orthotope. We compute representative values of these size measures of the elliptical sets. To determine reasonable size measures for elliptical regions, we proceed with an ellipse defined by
$E=\{x:x^{T}Kx\leq c\},$ (15)
where $K$ is an $m\times m$, p.d. matrix, and $c>0$. Since the center of the ellipse has no role in the size measures, we take it zero. Let $V_{E}(K,m,c)$ denote the Lebesgue volume of the set $E$. It can be easily calculated that
$V_{E}(K,m,c)=\frac{\pi^{m/2}}{\Gamma((m+2)/2)}\times\frac{c^{m/2}}{\mbox{det}(K)^{1/2}},$ (16)
where $\Gamma(\cdot)$ is the gamma function and $\mbox{det}(K)$ is the determinant of $K$. A standard length measure of an elliptical region is not available. We derive below a reasonable expression for it. For an $m$-dimensional rectangular region, its volume is the product of the volume of the $(m-1)$-dimensional of the cross-section of the set defined by the intersecting hyperplane, say, $\theta_{1}=d$, and the length $(U_{1}-L_{1})$ of original rectangular set along the $\theta_{1}$ direction (where $L_{1}<d<U_{1}$). Thus, the length of the set along the $x_{1}$ direction is the ratio of the volume of the original orthotope to the volume of the lower-dimensional cross-section orthotope. We will use the above discussion for the orthotope region as an analogy to come up with a suitable “length” for the ellipse along the $x_{1}$ direction. For the ellipse, we take for any $x_{1}$ and $x_{(2)}=(x_{2},\cdots,x_{m})^{T}$ with $(x_{1},x_{(2)}^{T})^{T}$ inside the set, we define a central cross-section by the hyperplane $x_{1}=-K_{12}x_{(2)}/K_{11}$, where we partition the matrix $K$ as $((K_{ij}))$, $i,j=1,2$, in conformity with the partitioning of $x$ above. We also define the p.d. matrix $K_{22.1}=K_{22}-K_{21}K_{11}^{-1}K_{12}$. The corresponding cross-section is $E_{C}=\{x_{(2)}^{T}K_{22.1}x_{(2)}\leq c\}.$ From equation (16), it similarly follows that
$V_{E_{C}}(K_{22.1},m-1,c)=\frac{\pi^{(m-1)/2}}{\Gamma((m+1)/2)}\times\frac{c^{(m-1)/2}}{\mbox{det}(K_{22.1})^{1/2}}.$ (17)
A reasonably representative length for the $x_{1}$ side of the ellipse is the ratio of the volume in (16) to the volume in (17). Simplification leads to an expression. From this, after simplification, in general, we have along the $x_{i}$ direction the representative length, $L_{R,i}$ (say), is
$L_{R,i}=B(\frac{1}{2},\frac{m+1}{2})\times\{\sqrt{c}/\sqrt{K_{ii}}\},~{}i=1,\cdots,m,$ (18)
where $B(a,b)$ is the regular beta function. The above proposal of the length of a side of an elliptical region is only heuristic. Indeed it is a reasonable requirement that these lengths should be “calibrated” by a suitable scale so that for the modified orthotope based on the lengths of the “scaled sides” will match the volume of the original ellipse. Based on this consideration, the modified lengths $L_{M,i}$ are given by
$L_{M,i}=\{V_{E}(K,m,c)/\prod_{i=1}^{m}L_{R,i}\}^{1/m}L_{R,i}.$ (19)
Mean of the $L_{M,i}$’s is a “linear” measure, $L_{E}$, of the elliptical set, and may be compared to $L_{C}$.

### v2

@@ 11 Summary and Conclusions
Table S.1: An extraction of a $9\times 9$ double stochastic matrix for the CD from the HB HPD (weighted) credible set for the commuting example
States
Rank ID KS IA AK MT NE WY ND SD
1 0.00 0.00 0.00 0.00 0.00 0.00 0.01 41.92 58.07
2 0.00 0.00 0.00 0.00 0.00 0.00 0.11 57.96 41.92
3 0.00 0.00 0.00 8.78 20.77 31.76 38.57 0.11 0.01
4 0.00 0.00 0.00 11.37 20.78 49.58 18.27 0.00 0.00
5 0.00 0.11 0.60 22.40 41.86 14.64 20.40 0.00 0.00
6 0.00 2.43 7.06 49.37 16.30 4.02 20.82 0.00 0.00
7 0.00 31.77 61.13 5.92 0.26 0.00 0.92 0.00 0.00
8 0.10 65.61 31.21 2.16 0.04 0.00 0.89 0.00 0.00
9 99.90 0.08 0.00 0.00 0.00 0.00 0.02 0.00 0.00
Table S.2: A table of average or expected absolute deviations of ranking from the surrogate median income data from 1990 census
@@
S.4 Ranking of baseball players: An example by Efron and Morris (1975)
Efron and Morris (1975) considered batting averages of 18 major league baseball batters from their first 45 at bats in the 1970 baseball season to predict their performances in the remainder of that season. Let $\theta_{i}$ be the unknown true batting average of the $i$th player during that season. A reasonable representative value for $\theta_{i}$ was the player’s average ($\omega_{i}$) in the remainder of the season after his first 45 at bats. The sample proportion of hits $Y_{i}$ from first 45 at bats is approximately normal with mean $\theta_{i}$ and variance $D_{i}$. In our analysis, we use the estimated variance, given by $D_{i}=Y_{i}(1-Y_{i})/45.$ Treating the available $\omega_{i}$ as the “gold standard” value of $\theta_{i}$, the total empirical squared error (TESE), $\sum_{i=1}^{18}(y_{i}-\omega_{i})^{2}$, for the direct estimates $y_{i}$’s, is $0.07537$, and the same error for the HB and the empirical best linear unbiased predictor (EBLUP) estimates of $\theta_{i}$’s are $0.02368$ and $0.02356$, respectively. The direct estimates have $3.18$ times larger TESE than the HB estimates, and $3.20$ times larger TESE than the EBLUP estimates.
Efron and Morris (1975) used the arc-sin transformed parameters $\psi_{i}=arcsin(2\theta_{i}-1)$, for the transformed data $\hat{\psi}_{i}=arcsin(2Y_{i}-1)$ in their EB analysis. This transformation resulted in equal approximate sampling variances, $1/45$, for all 18 $\hat{\psi}_{i}$’s. They applied Stein’s shrinkage estimation method based on an exchangeable model to calculate ${\hat{\psi}}_{J-S,i}$, the James-Stein estimate of $\psi_{i}$. The TESE of prediction for the direct estimates $\hat{\psi}_{i}$ from $arcsin(2\omega_{i}-1)$ is 0.3902, which is about 3.5 times of 0.1113, the TESE for the James-Stein estimates ${\hat{\psi}}_{J-S,i}$. This application showed a substantial efficiency of Stein’s method. The efficiency for HB estimators of $\psi_{i}$ based on the same model is about $3.36$, which is comparable. Note that these numbers are comparable to $3.18$ and $3.20$, we obtained above for EBLUP and HB predictors of $\theta_{i}$’s without transformation.
Figure S.2: An alternative visualization of the Klein et al. (2020) 90% joint confidence region for overall ranking of baseball players with overlapping CDs for UB and HB models (weighted, elliptical method)
This confidence set for the rank vector is depicted in Figure S.2. Possible ranks for each player in this set are shown by a yellow line segment stretching from 1 to 18. According to this confidence set, any player can rank from 1 to 18, and any particular rank can be held by any of these 18 players; clearly, it is totally uninformative. Due to multiple testing, the critical value $2.76$ (cf. equation (13) of Klein et al. (2020)) is too large to declare any comparison between two direct estimates to be significant. A joint $90\%$ credible distribution for the rank vector is a double-stochastic matrix, each row and each column of the matrix sums to one. This probability matrix is pictorially presented by overlaying on the figure of the confidence set solid red circles with area of a circle is proportional to the probability it is representing. The top matrix is based on the elliptic set for $\theta$ under the UB model and the bottom matrix is a similar version for the intercept only HB model. From the figure, it is clear that many of the probabilities in this matrix are practically zeros implying that those combinations of players and ranks are improbable. Undoubtedly, the credible distributions are lot more informative than the confidence set.
The observed rank $R_{i}$ of the $i$th player based on $y$ is shown by a blue square on the $R_{i}$th row and the $i$th column of the figure. If several players tie, we follow the recommendation of Klein et al. (2020) to assign the highest possible rank to the tied players. We used $\omega_{i}$ values to find the surrogate rank vector $\xi$ for the true vector $\check{r}$. The $i$th component, $\xi_{i}$, for player $i$ is similarly depicted by a black circle on the $(\xi_{i},i)$ cell of the matrix.
The approach to the solution in the top panel of Figure S.2 is the one which is the most similar to the Klein et al. (2020) solution. It is clear from the first row of this panel is that Player 1 has the largest probability of being the best hitter, and only the first five or six players have realistic chances to be the best hitter. Similarly, some of the players 12 through 18 are the most likely ones who had the lowest performance. These lists of players are much narrower than the whole set of $18$ players suggested by Klein et al. (2020) method. From the first column of this panel, on the other hand, we find that a realistic list of ranks for Player 1 is $\{14,15,16,17,18\}$. This list is much narrower than the entire set $\{1,2,\cdots,18\}$ suggested by the frequentist method. This figure also indicates that most likely ranks for player 17 are 1 to 7. The rank $17$ is most likely to be held by players 1 to 7. Clearly, the Bayesian solution is substantially more informative.
In the bottom panel of Figure S.2 we depicted the double stochastic matrix from the HPD credible set for $\theta$ using the HB model. Due to shrinkage of the joint posterior distribution towards a vector of common means, probability distributions of the rankings of the few uppermost and the lowermost players are not as sharp as those provided by the corresponding credible distributions from the UB model. Even though there is a weakness of the HB method for its reliance on shrinkage to a common mean in the absence of a useful covariate, the method provides better solutions than the UB method for some players whose early batting averages are at odds with their latter scores. Here, five players (4, 8, 13, 16, 17) showed a large or moderately large discordance between their early and late season performances. Two most discordant players are 4 and 17. Player 4 did not perform well in the remainder of the season which earned him a very low surrogate rank $3$ compared to his high $y$ rank $15$. From the top panel of Figure S.2 we see that whereas the UB method assigns, perhaps erroneously, nearly 87% probability of Player 4 having a rank higher than 9, the bottom panel of Figure S.2 reveals that the HB method correctly assigns less than 63% probability to this player having such a high rank. At the other end, Player 17 performed very well in the remainder of the season which earned him a very high surrogate rank $17$ (vs. a very low $y$-rank 2). Here, again, Figure S.2 for the unweighted UB Cartesian method assigns, perhaps erroneously, nearly 93% probability of it having a rank lower than 10, the bottom panel of Figure S.2 for the HB method correctly assigns less than 67% probability to it having such a low rank. Revised rankings of the discordant players afforded by the HB method are noteworthy and more reliable. We have demonstrated through simulation in Subsection 10.1 and the median income example in Section 9 that the HB method in the presence of useful covariates is substantially more effective.
Also, for each player the variance of the possible ranks in the joint Klein et al. (2020) confidence region is $26.9$, the largest possible value, and the corresponding posterior variances from the credible distributions of ranks for these players vary between 5.36 and 16.36 for the UB model, and between 16.75 and 23.40 for the HB model. These numbers show a superiority of the credible distributions, created from the UB model or the HB model. We provide additional comparisons of the frequentist and the Bayesian solutions in Table  S.3.
Table S.3: Batting averages of 18 Major League players in 1970: $y_{i}$ is the average from first 45 at bats, $\omega_{i}$ the remainder of the season average, $R_{i}$ rank of $y_{i}$, $\xi_{i}$ rank of $\omega_{i}$, and deviation measures described in equation (14) for five overall ranking set estimators. For each of the two Bayesian models, we have two methods based on elliptical (subscript E) and Cartesian (subscript C) sets. We used the weighted method here.
1. Clemente .400 .346 18 18 5.90 5.90 1.18 1.18 8.50
2. F. Robinson .378 .298 17 15 4.39 4.39 1.87 1.87 6.17
3. F. Howard .356 .276 16 13 4.24 21.0 2.73 2.73 5.17
4. Johnstone .333 .222 15 3 8.11 8.11 11.40 11.39 6.83
5. Berry .311 .273 13.5 12 4.33 4.33 2.44 2.44 4.83
6. Spencer .311 .270 13.5 11 4.25 4.25 2.96 2.96 4.61
7. Kessinger .289 .263 12 7 4.96 4.96 4.92 4.92 4.83
8. L. Alvarado .267 .210 11 2 7.66 7.66 8.13 8.13 7.61
9. Santo .244 .269 9.5 10 4.36 4.34 3.09 3.09 4.50
10. Swoboda .244 .230 9.5 5 5.41 5.41 3.98 3.98 5.61
11. Unser .222 .264 6 8.5 4.26 4.26 3.09 3.08 4.56
12. Williams .222 .256 6 6 4.40 4.40 2.64 2.64 5.17
13. Scott .222 .303 6 16 7.64 7.64 9.36 9.36 6.83
14. Petrocelli .222 .264 6 8.5 4.25 4.25 3.33 3.33 4.56
15. E. Rodriguez .222 .226 6 4 5.38 5.38 3.46 3.46 6.17
16. Campenaris .200 .285 3 14 6.64 6.64 8.97 8.97 5.61
17. Munson .178 .316 2 17 9.54 9.54 13.57 13.56 7.61
18. Alvis .156 .200 1 1 5.46 5.46 1.43 1.43 8.50
Total abs. dev. ($\check{r}$) 101.16 101.16 88.56 88.56 107.64
$10^{13}\times$ Vol. ($\theta$) 51.38 1.32 93756 1757 85722
Average length ($\theta$) 0.236 0.193 0.359 0.288 0.357
From Table S.3 we find that the Klein et al. (2020) method has the highest sum of average absolute deviations (based on equation (14), about 6.4% larger than for the HB method, and about 21.5% larger than for the UB method. This is counter-intuitive as the shrinkage estimators of $\theta_{i}$’s from the exchangeable HB model have lower total squared error than the same for the direct estimators $Y_{i}$’s, which are Bayes estimates of $\theta_{i}$ under the UB model. We have noted above that the total squared error of UB estimate of $\theta$ from the “true” vector $\omega$ was $0.07537$, and for the HB estimate it was $0.02368$. We present below a possible explanation for the domination of the UB method over the HB method when it comes to inference for overall ranking based on their credible distributions.
In this example, estimating $\theta_{i}$ from the HB model shrinks the direct estimate; the shrinkage of $y_{i}$ towards a grand mean, measured roughly by the posterior mean of $E[D_{i}/(D_{i}+A)|y]$, is substantial (ranging between $0.75$ and $0.63$, and the shrinkage is more for the larger $D_{i}$’s that correspond to larger $y_{i}$’s. In our example, $D_{i}$’s decrease from $0.0053$ to $0.0029$, and the posterior mean of $A$ is $0.0024$ and the median of $A$ is $0.0017$.) These indicate that even in comparison with the small $D_{i}$’s, $A$ is small. In the absence of a covariate, a very small model error variance for an intercept only model essentially forces that $\theta_{i}$’s are all nearly equal. Accurate inference for the rank vector for such $\theta$ will naturally be challenging. Among the components of $\theta$, any differentiating information about $\theta_{i}$’s is given by its individual $y_{i}$’s only. Clearly, any shrinkage of an estimator of $\theta_{i}$ to a common value will reduce the estimator’s reliance on individual $y_{i}$’s, and thus the HB posterior distribution of $\theta_{i}$’s for ranking will be less informative than $y_{i}$, which is the center of the individual UB posterior distributions.
Due to lack of a covariate, corresponding to the set of larger observations (the first 4 or 5 observations which are in descending order), the HB posterior distribution of any of these $\theta_{i}$’s will shrink downward (to some grand mean). This downward shrinking of the posterior distribution of $\theta_{i}$’s will lower the probability of them holding the higher ranks (18 through 15 or 14). In fact, from Table S.4 the posterior expectation of $\check{r}_{1}$ in our example is $12.1$, which is quite a bit smaller than the largest rank $18$ (in this case) it is supposedly estimating. Similarly, for the next two smaller observations, the posterior expected ranks are again quite smaller than the target ranks 17 and 16. These discrepancies contribute to a bigger absolute deviations reported in Table S.3. On the other hand, the distributions of $\check{r}_{1},\check{r}_{2},\check{r}_{3}$ for the UB model do not shrink downward much since the posterior of $\theta_{i}$’s are independent and each is concentrated around its $y_{i}$. The corresponding posterior means of $\check{r}_{1},\check{r}_{2},\check{r}_{3}$ are $16.8,16.2$ and $15.2$ (cf. Table S.4), which are not too far from $18,17$ and $16$, respectively. These lesser discrepancies make smaller contributions to the expected absolute deviations reported Table S.3. A similar shrinkage upward will be noticed for the set of smaller observations (observations 14 through 18). Corresponding expected absolute deviations of the rank from the surrogate true rank $\xi_{i}$ for observations 14 through 18 under the HB model again tend to be bigger than their counterparts from the UB model. We see bigger discrepancies between the HB ranks and the UB ranks for the first 4 and the last 4 observations. All of these mostly explain the smaller total posterior expected absolute rank deviations for the UB model in comparison with their counterparts for the HB model.
Table S.4: Estimation of ranks of baseball players by different methods (weighted versions were used for Bayesian methods)
Players P01 P02 P03 P04 P05 P06 P07 P08 P09
$y$ ranks 18 17 16 15 14 13 12 11 10
$\omega$ ranks 18 15 13 3 12 11 7 2 10
KWW 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50
HB-Cart 12.10 11.86 11.60 10.94 10.41 10.94 10.27 9.57 9.35
HB-Ellip 12.10 11.86 11.60 10.94 10.41 10.94 10.27 9.57 9.35
UB-Cart 16.82 16.20 15.23 14.40 12.98 12.80 11.51 10.12 8.40
UB-Ellip 16.82 16.20 15.23 14.39 12.98 12.79 11.51 10.12 8.40
Players P10 P11 P12 P13 P14 P15 P16 P17 P18 Average
$y$ ranks 9 8 7 6 5 4 3 2 1 9.5
$\omega$ ranks 5 8.5 6 16 8.5 4 14 17 1 9.5
KWW 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.50 9.5
HB-Cart 9.22 8.61 8.54 8.60 8.56 8.56 7.91 7.50 6.46 9.5
HB-Ellip 9.22 8.61 8.54 8.60 8.56 8.56 7.92 7.50 6.46 9.5
UB-Cart 8.41 6.71 6.69 6.64 6.43 6.79 5.03 3.43 2.43 9.5
UB-Ellip 8.41 6.72 6.69 6.64 6.43 6.79 5.03 3.44 2.43 9.5
Even though the UB method is poised to have smaller total expected absolute deviations than the HB method without a covariate, for some individual players the HB method is better suited to estimate ranks than the UB method. Continuing the above discussion further, since the posterior distribution of $\theta_{i}$ under the UB model is determined only by $y_{i}$, in case of a big discrepancy between $y_{i}$ and the unknown $\theta_{i}$ it is estimating (which will be reflected through a bigger difference between $y_{i}$ and $\omega_{i}$, the surrogate value of $\theta_{i}$), the posterior expected absolute deviation of $\check{r}_{i}$ from $\xi_{i}$ will be large ($\check{r}_{i}$, the rank of $\theta_{i}$, and $\xi_{i}$, the rank of gold standard value $\omega_{i}$). Indeed this happens for five players 4, 8, 13, 16 and 17 with large discordance between a player’s early and late season performances. Among these five players, player 17 has the second smallest observation, who displayed large discrepancies between $y_{i}$ and $\omega_{i}$. Deviations $\epsilon_{i,UB}$ for these players are much larger than the average deviations of the other players, and these deviations also exceed the corresponding players’ deviations $\epsilon_{i,HB}$ based on the exchangeable HB model. If one or more player substantially over-perform or under-perform in the remainder of the reason, based on their later performances the true ranks of these players will change substantially from their ranks determined by $y_{i}$’s. The HB method by virtue of its borrowing from the other players is better suited to rectify the erroneous ranks of the players based on $y_{i}$’s. This is the likely reason that the HB ranks for the five players identified above perform better than the UB ranks that are based on individual $y_{i}$ alone; a downward shrinkage of the HB posterior distribution of the over-performing (in their first 45 at bats) players 4 and 8, and an upward shrinkage of the HB posterior distributions of the under-performing (in their first 45 at bats) players 13, 16 and 17. The HB posterior distributions of $\theta_{i}$ of these players, by virtue of their borrowing information from the other players, are substantially pulled toward the “true value”, measured by $\omega_{i}$. This is a useful property of the HB model. No such borrowing of information is possible through the UB model as the posterior distributions of the components $\theta_{i}$’s are independent and centered around $y_{i}$.
Lastly, from this example where the total posterior expectation of absolute deviations of the true rank $\check{r}_{i}$ from the surrogate rank $\xi_{i}$ for the UB model is lower than the same for the HB model, we may not conclude that the UB model will always provide a more accurate solution. In fact, it is shown in simulations in the paper that if a covariate effectively explains the $\theta_{i}$’s, the credible distributions of overall ranking based on the HB model are superior to the credible distributions derived from the UB model. Intuitively, this occurs since the posterior distribution of $\check{r}_{i}$ under the HB model will better concentrate near the true rank than the same posterior distribution under the UB model. We should note that even though in terms of total of expected absolute deviations, the frequentist Klein et al. (2020) method fares the worst, it is surprisingly robust for the five under-/over-performing players 4, 8, 13, 16 and 17. Finally, we included some size measures of the confidence set (for the KWW method) and credible sets for Bayesian methods, the sets are critical for the ranking problem. The last two rows of Table S.3 provide volumes and approximate average lengths of the sets, respectively. In terms of volume, the HB sets are at least a thousand times better than the other sets. In terms of length, the HB sets are at least 50% better than the other sets. The best among all five sets is the HB elliptical set.